# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

### v2

@@ 5 Experimental Setup
Models.
We use resnet18 [34] and vit_b_16 [35] as FMs, one CNN-based and one Transformer-based architecture. In neural collapse investigations (section 6.3), we also work with resnet50 [34] and swin_b [36]. All four models are pre-trained on ImageNet-1K (ILSVRC 2012) [37]. To generate the embeddings, we use PyTorch’s torchvision models.
Tasks.
To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10 [40], each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
In the investigations of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet). As its test set is not available, we use its validation set in $\mathcal{NC}_{1}$ computations. For ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis. These tasks are called "$d$ in $1$", $d=1,\dotsc,10$, each containing $10$ different labels. More precisely, we take 100 random classes from ImageNet and sequentially map the first $d$ to label $1$, the second $d$ to label $2$, etc., until we reach 10 distinct labels. See fig. 4 for a simplified illustration. We do this random sampling 10 times which results in $10\cdot 10=100$ tasks. Furthermore, to ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
Figure 4: Combining multiple classes into one to create tasks with multi-modal class distributions. Example: "$d$ in 1", $d=2,3$.
@@
In total, we run approximately 150 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds. This is feasible with minimal effort as imprinting is a highly efficient method that operates without relying on gradient descent or other non-linear optimization techniques.
@@
Throughout our experiments, the median accuracy on the test set for three different seeds is reported, if not otherwise specified. In sections 6.1 and 6.2, we investigate the imprinting performance by varying the FM (2) and T (12). We then sort the combinations by their final accuracy. There are $2\cdot 12=24$ potentially different ranks for each of the combinations. We show the average rank, average accuracy, and statistical significance in ranking (dis-)agreements through critical difference (CD) diagrams as presented in [41]. [^2] The code used to generate these diagrams is inspired by [42]. In the CD diagrams, a thick horizontal line indicates a group of combinations that are not significantly different from each other in terms of accuracy. We consider differences significant if $p<0.05$.
In experiments with neural collapse (section 6.3), we investigate four FMs on 100 ImageNet tasks and the three tasks containing all of MNIST, FashionMNIST, and CIFAR-10, respectively.
@@ 6 Results
Our main experimental insights are:
  1. 1.
    Our imprinting framework generalizes previous methods, and we find a new superior imprinting strategy (section 6.1).
  2. 2.
    We show that our strategy is beneficial in few-shot scenarios with as little as 50 samples per class (section 6.2).
  3. 3.
    We identify a correlation between imprinting success utilizing multiple proxies and measures of neural collapse (section 6.3).
@@ 6.1 Best Imprinting Strategy
We provide a comparison between memory-constrained methods used for imprinting on foundation models in fig. 2, namely, [4, 14, 15], as well as a novel configuration that results from our framework which we call “Ours”. We investigate the impact of using m-nn aggregation on all data afterward. We focus on $k=20$ and find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all other approaches by a margin of $4\%$ with statistical significance, as can be seen in the CD diagram. Below, we analyze each of the components of the framework separately.
@@
To analyze the impact of GEN, we first focus on the max aggregation. For the weight generation analysis, we do not fix NORM, but simply show the run with the best NORM combination, if not otherwise specified. The m-nn aggregation and different values for NORM are analyzed later in this section.
GEN $k$ Avg. acc. %
k-means 20 91.44
k-medoids 20 87.31
mean 1 87.76
k-cov-max 20 84.74
k-random 20 82.86
k-fps 20 66.23
@@
Figure 6: Benchmarking different GEN methods with resnet18 on CIFAR-10. All combinations employ L2 for all NORM.
Initially, we limit the number of generated proxies ($k\leq 20$). Results in fig. 5 show how k-means, using as many proxies as possible (in this case, $20$) outperforms by $4\%$ on average accuracy compared to all the other GEN methods. The CD diagram illustrates its statistical significance in ranking. Furthermore, while k-medoids with $20$ proxies is computationally expensive, it is statistically on par with mean, and covariance maximization, furthest-point sampling and random selection show even weaker performances. We find similar results for $k\leq 5$, where k-means outperforms the other methods as well (see fig. 13).
As the number of proxies ($k$) increases, k-means continues to be the best GEN method. An example for resnet18 and CIFAR-10 can be found in fig. 6. All methods converge towards the point of imprinting (saving) all data ($k=\text{all}$), even surpassing it in the case of k-means. Due to its superior performance, we mainly focus on k-means in the remainder of the analysis.
Normalization (NORM).
To investigate the role of normalization, we compare all the different NORM methods, focusing on k-means as GEN with $k=1,5,20$. For $k=1$ and varying NORMpost (while taking best values for NORMpre and NORMinf implicitly), fig. 7 shows that for weight normalization, L2 is by far the best choice. quantile and none normalization perform significantly worse.
NORMpost Avg. acc. %
L2 87.76
quantile 80.70
none 84.04
@@
Figure 7: Benchmarking NORMpost mechanism across FMs and Ts. The best NORMpre and NORMinf combinations for each row are used implicitly. GEN is fixed to mean (that is, $k=1$) and AGG is fixed to max. The CD diagram shows the statistical significance of L2 as weight normalization NORMpost.
NORMpre NORMinf Avg. acc.
none L2 87.76
none none 87.76
L2 L2 87.72
Keeping L2 for NORMpost fixed, we find no statistical differences between the different combinations of NORMpre and NORMinf. Its performances can be found in table 1. For larger values of $k$, the differences among NORMpost become even more pronounced, but for NORMpre and NORMinf, it stays statistically indifferent for L2 weight normalization (see tables 2 and 14 for all combinations at once with $k=1$, and tables 3 and 15 for $k=20$).
Henceforth, we limit all the succeeding experiments to Qi’s [4] normalization, that is, using L2 for all NORM. We choose this combination of normalizations to specifically capture cosine similarity in max aggregation.
@@
In addition to max, we study the effect of m-nn as an aggregation method. Recall that max is a special case of m-nn when $m=1$ (as NORMpost is set to L2). We investigate different values for $m\in\{1,3,5,20,50\}$.
AGG Avg. acc. %
5-nn 93.86
3-nn 93.68
20-nn 93.73
1-nn 92.97
max 92.97
50-nn 93.15
[…]
Figure 8: Benchmarking AGG mechanism across FMs and Ts. GEN is fixed to all ($k=\text{all}$), that is, imprinting (saving) all data to weights. L2 normalization is used for all NORM. The CD diagram shows statistical significance of 3-nn, 5-nn, and 20-nn over max aggregation.
When all data is imprinted, fig. 8 shows that using m-nn aggregation for $m\in\{3,5,20\}$ is slightly better than max.
AGG Avg. acc. %
1-nn 91.48
max 91.48
3-nn 91.01
5-nn 90.61
20-nn 87.64
[…]
With $k=20$ and k-means as GEN, max ($=$1-nn) aggregation becomes the top performing combination (see fig. 9). Furthermore, the reduction of proxies (from all $\approx 6000$ per class to $k=20$) leads to a minor decrease in accuracy of around $2\%$.
6.2 Few-Shot Scenario
We analyze the n-shot scenario with our method, that is, k-means as GEN, L2 normalizations, and max as AGG. Furthermore, we focus on the large tasks T, that contain ten classes at once. In this scenario, due to sampling only a few examples ($n$), we average over five different seeds.
Figure 10: k-means with different values for $k$ in n-shot scenarios. $95\%$ confidence intervals are shown in shaded colors. Other variables are fixed to our previously described best method of using L2 normalizations and max as AGG. Note that only data for the meaningful case of $k\leq n$ is shown. It can be inferred that for MNIST and FashionMNIST, mean is not the best strategy anymore starting at roughly 50 samples.
From the results shown in fig. 10, we find that as the number of samples $n$ increases, k-means starts outperforming mean imprinting. The usage of a higher number of proxies $k$ results in even greater performance. This shift occurs at roughly 50 samples per class for MNIST and FashionMNIST, while for CIFAR-10, $k>1$ becomes prominently better at around 200 samples per class (see fig. 16 for a focus on $10\leq n\leq 400$).
@@ 6.3 Neural Collapse and Number of Proxies
Figure 11 depicts the neural collapse measurement $\mathcal{NC}_{1}$ (see eq. 1) for the three tasks containing all of MNIST, FashionMNIST, CIFAR-10, as well as the 100 ImageNet tasks with remapped labels as explained in section 5. We can see that ImageNet has a close-to-zero $\mathcal{NC}_{1}$ score, which increases linearly when adding more classes to each label (i.e., increasing multi-modality). As for other datasets, CIFAR-10 is generally more collapsed according to its low value of $\mathcal{NC}_{1}$. We hypothesize that this is due to the similarity of its categories to those appearing in ImageNet. Apart from that, the $\mathcal{NC}_{1}$ for Transformer-based architectures is much lower and therefore they are more collapsed compared to the CNN-inspired FMs. Architectural differences are further explained in section A.2.
Figure 11: Average $\mathcal{NC}_{1}$ and $95\%$ confidence intervals of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$. The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted as dots. When compared to $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures. A clear linear relationship across $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
For the same data, fig. 12 depicts accuracy over a varying number of proxies $k$ inferred from k-means. A prominent peak at $k=d$ can be inferred for every FM, and reflects that $d$ class proxies lead to the best result for $d$-modal class distributions. The fact that CIFAR-10 has the lowest $\mathcal{NC}_{1}$ (see fig. 11) is reflected by flat curves over $k$. This confirms that the $\mathcal{NC}_{1}$ score is a significant indicator of multi-modality. Namely, a higher $\mathcal{NC}_{1}$ score indicates the benefits of using a higher number of proxies.
Furthermore, increasing $k$ for the ImageNet sets has a much larger effect on the CNN-based models. We argue that this is because of their higher values of $\mathcal{NC}_{1}$ and investigate this more deeply in section A.2.
@@
Figure 12: Averaged accuracy of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$ over number of proxies $k$ used for k-means in GEN. $95\%$ confidence intervals are shown in shaded colors. L2 for all NORM and max as aggregation are used. Accuracies of the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are shown in dotted lines. In all four plots, peaks in accuracy at $k=d$ can be inferred. This confirms the connection between the effect of using multiple proxies and the collapse of the data.
7 Discussion

### v3

@@ 3.2 Quantifying Neural Collapse
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$) as in zhu2021geometric. Given a foundation model FM and a finite dataset containing $C$ classes with (for simplicity) $N$ samples per class, we use its $l$-dimensional embeddings $\{\boldsymbol{\mathbf{h}}_{c,i}\}_{1\leq c\leq C,\;1\leq i\leq N}$ given by FM to define the global feature mean $\boldsymbol{\mathbf{h}}_{G}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, class means $\overline{\boldsymbol{\mathbf{h}}}_{c}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, the within-class covariance matrix $\mathbf{\Sigma}_{W}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)^{\top}$, and the between-class covariance matrix $\mathbf{\Sigma}_{B}\coloneqq\frac{1}{C}\sum_{c=1}^{C}\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)^{\top}$ to finally compute
[…]
where + symbolizes the pseudo-inverse.
Based on eq.˜1, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse. In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig.˜11(a)). Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $C$. According to NC, imprinting the mean, as originally done by qi-imprinting, is best when $\mathcal{NC}_{1}$ is small. We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class. We investigate this in section˜5.3.
3.3 Significance Testing with Critical Difference Diagrams
Our experiments compare large numbers of imprinting configurations across tasks with substantially different achievable accuracy levels. To obtain scale-independent statistical comparisons, we base our analysis on ranking (dis-)agreements rather than raw accuracy. Non-parametric rank-based tests are well suited for this setting. Following cddiag, we use the Friedman test and Wilcoxon signed-rank tests because they do not assume normality or homoscedasticity, and can be applied uniformly to any evaluation metric. Empirical evidence in cddiag further suggests that these tests have good power in classifier comparison scenarios.
Concretely, for each evaluation instance (i.e., foundation model FM and task T) we compute the classification accuracy of all configurations and assign ranks (rank 1 = highest accuracy). This yields a matrix of ranks with one row per configuration and one column per evaluation instance. On this matrix we first apply the Friedman test to assess whether there are any overall differences between configurations. If the Friedman test rejects the global null hypothesis at significance level $\alpha=0.05$, we proceed with a post-hoc, pairwise analysis. Specifically, for every pair of configurations we run a two-sided Wilcoxon signed-rank test across evaluation instances. The resulting $p$-values are corrected for multiple comparisons using the Holm step-down procedure (holm1979simple).
Critical difference (CD) diagrams summarize both the average performance and the significance structure. The horizontal axis shows the average rank of each configuration across all evaluation instances, with lower ranks indicating better performance. Two configurations are connected by a thick horizontal line if their Wilcoxon–Holm comparison does not show a significant difference.
@@ 4 Experimental Setup
Foundation Models FM.
We use resnet18, resnet50 (resnet), vit_b_16 (vit), and swin_b (swin) as FMs, two CNN-based and two Transformer-based architectures. All four models are pre-trained on ImageNet-1K (ILSVRC 2012) (imagenet). To generate the embeddings, we use PyTorch’s torchvision models.
Tasks T.
We analyze multi-class classification scenarios without separating base and new classes, instead focusing on all classes within a novel T at the same time. To investigate the effect of the number of samples given, we look at n-shot ($n\in\mathbb{N}$) scenarios. For that, we randomly pre-sample the training data of T to $n$ samples per class – transitioning into the low-data regime.
To find out the best imprinting strategy within our IMPRINT framework, we focus on tasks T created from the datasets MNIST (mnist), FashionMNIST (fashionmnist), and CIFAR-10 (cifar10), each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
Neural Collapse.
In the analysis of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet). As its test set is not available, we use its validation set in $\mathcal{NC}_{1}$ computations. Furthermore, for ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis. These tasks are called $d$ in 11, $d=1,\dotsc,10$, each containing $10$ different labels. More precisely, we take 100 random classes from ImageNet and sequentially map the first $d$ to label $1$, the second $d$ to label $2$, etc., until we reach 10 distinct labels. See fig.˜4 for a simplified illustration.
Moreover, we construct a new out-of-distribution, non-collapsed dataset by merging all classes from four digit datasets: MNIST, MNIST-M (mnistm), SVHN (svhn), and USPS (usps). The resulting dataset, which we refer to as CombiDigits, exhibits reduced collapse due to the greater distributional diversity within each class. To ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
@@
In total, we ran approximately 500 000 experiments, varying imprinting components, foundation models, tasks, and seeds. This is feasible with minimal effort as imprinting is a highly efficient method: core steps such as weight generation (GEN), normalization (NORM), and aggregation (AGG) are linear in dataset size or number of proxies and parallelize efficiently within each step. Details on experimental infrastructure and parallelization are provided in section˜A.6.
Figure 4: Combining multiple classes into one to create tasks with multi-modal class distributions. Simplified example for $d$ in 1”, $d=1,2,3$, with only six (instead of 100) original classes.
@@
Unless stated otherwise, we report the median test accuracy over three different seeds. In sections˜5.1 and 5.2, we evaluate the imprinting performance by varying the FM (4) and T (12) and report average accuracy and standard deviation (std) across these. Due to the heterogeneity of models and tasks, large std are expected. Therefore, overall method comparisons are based on ranks rather than absolute accuracies. Methods are ranked by their median accuracy, yielding $4\cdot 12=48$ potentially different ranks. We report the average rank and assess statistical significance of ranking (dis-)agreements using critical difference (CD) diagrams as explained in section˜3.3. The code used to generate these diagrams is inspired by cddiag-code.
In experiments with neural collapse (section˜5.3), we investigate the four FMs on 100 random ImageNet tasks with remapped labels (as explained above) and the four tasks containing all of MNIST, FashionMNIST, CIFAR-10, and CombiDigits, respectively.
@@ 5 Results
Our main experimental insights are: 1. Our IMPRINT framework generalizes previous methods, and we find a new superior imprinting strategy (section˜5.1). 2. We show that our novel strategy is even beneficial in low-data regimes with as little as 50 samples per class (section˜5.2). 3. We identify a correlation between imprinting success utilizing multiple proxies and a measure of neural collapse (section˜5.3).
@@ 5.1 Best Imprinting Strategy
We provide a comparison between existing memory-constrained methods used for imprinting on foundation models in fig.˜2, namely, the work by qi-imprinting; itsdone; ncmbaseline, as well as a novel, best-performing configuration (“Ours”) that results from IMPRINT. Focusing on $k=20$, we find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all previously studied approaches by a margin of $4\%$ on average with statistical significance. For reference, we additionally report an oracle baseline that uses cross-class feature statistics to generate its weights, representing an upper bound that is not constrained by imprinting. The results indicate that our new imprinting method significantly narrows the gap between single-proxy mean imprinting and this oracle baseline.
The impact of using the unconstrained m-nn aggregation on all data is investigated at the end of this section. Next, we analyze each of the components of IMPRINT separately.
[…]
[table]A table beside a figure GEN $k$ Avg. acc. % $\pm$ std k-means 20 91.04 $\pm 6.26$ k-medoids 20 87.01 $\pm 8.23$ mean 1 86.84 $\pm 7.80$ k-cov-max 20 83.98 $\pm 9.56$ k-random 20 82.14 $\pm 10.15$ k-fps 20 65.56 $\pm 12.32$ Refer to caption
@@
To assess the impact of GEN, we first focus on the max aggregation and do not fix NORM, but simply show the run with the best NORM combination, if not otherwise specified. The m-nn aggregation and different values for NORM are analyzed later in this section.
Initially, we limit the number of generated proxies to $k\leq 20$. Results in fig.˜5 show how k-means, using as many proxies as possible (in this case, $20$) outperforms by $4\%$ on average accuracy compared to all the other GEN methods. The CD diagram illustrates its statistical significance in ranking. Furthermore, while k-medoids with $20$ proxies (which necessarily have to be among the given samples, see section˜3.1) is computationally expensive, it is on par with mean, and covariance maximization, furthest-point sampling and random selection show even weaker performances. We find similar results for $k\leq 5$, where k-means outperforms the other methods as well (see fig.˜A.1).
Figure 6: Benchmarking different GEN methods with resnet18 on all of CIFAR-10 shows superiority of k-means proxies. All combinations employ L2 for all NORM and max as AGG.
As the number of proxies $k$ increases, k-means continues to be the best GEN method. An example for resnet18 and CIFAR-10 can be found in fig.˜6. All methods converge towards the point of imprinting (saving) all data ($k=\text{all}$), even surpassing it in the case of k-means. Due to its superior performance, we mainly focus on k-means in the remainder of the analysis.
Learned Weights.
Employing gradient-based methods – such as setting weights via the analytical least-squares initialization proposed by harun2025good, which uses data and labels from all classes jointly – can be seen as an upper bound (“Oracle” in fig.˜2) for the unsupervised weight imprinting approaches discussed here. We extend this method by combining it with k-means, forming k-least-squares, and observe in section˜A.5 that using multiple proxies per class instead of a single weight vector can still improve performance, particularly on datasets with high $\mathcal{NC}_{1}$.
Normalization (NORM).
We compare all the different NORM methods, focusing on k-means as GEN. For $k=1$ and varying NORMpost (while taking best values for NORMpre and NORMinf implicitly), fig.˜7 shows that L2 is the best choice for weight normalization. quantile and none normalization both perform worse.
@@
[table]A table beside a figure NORMpost Avg. acc. % $\pm$ std L2 86.84 $\pm 7.80$ quantile 82.90 $\pm 12.87$ none 83.26 $\pm 9.18$ Refer to caption
Figure 7: Benchmarking NORMpost mechanism across FMs and Ts. The best NORMpre and NORMinf combinations for each row are used implicitly. GEN is fixed to mean (that is, $k=1$) and AGG is fixed to max. The CD diagram shows the statistical significance of L2 as the best weight normalization NORMpost.
NORMpre NORMinf Avg. acc. % $\pm$ std
none L2 86.84 $\pm 7.80$
none none 86.84 $\pm 7.80$
L2 L2 86.79 $\pm 7.83$
Keeping L2 for NORMpost fixed, we find no statistical differences between the different combinations of NORMpre and NORMinf (see table˜1). For larger values of $k$, the differences among NORMpost become even more pronounced. However, the performance of NORMpre and NORMinf remains statistically indifferent for L2 weight normalization (see fig.˜A.2 for all combinations at once with $k=1$, and fig.˜A.3 for $k=20$).
We restrict all the subsequent experiments to using L2 across all NORM following qi-imprinting. This combination of normalizations is chosen to specifically model cosine similarity within max aggregation.
@@
[table]A table beside a figure AGG Avg. acc. % $\pm$ std 5-$n$ 93.74 $\pm 4.97$ 3-$n$ 93.50 $\pm 5.22$ 20-$n$ 93.56 $\pm 4.93$ 1-$n$ 92.81 $\pm 5.69$ max 92.81 $\pm 5.69$ 50-$n$ 92.91 $\pm 5.25$ Refer to caption
Figure 8: Benchmarking AGG mechanism across FMs and Ts. GEN is fixed to all ($k=\text{all}$), that is, imprinting (saving) all data to weights. L2 normalization is used for all NORM. The CD diagram shows statistical significance of 3-$n$, 5-$n$, and 20-$n$ over max aggregation.
[…]
[table]A table beside a figure AGG Avg. acc. % $\pm$ std 1-$n$ 91.06 $\pm 6.21$ max 91.06 $\pm 6.21$ 3-$n$ 90.59 $\pm 6.16$ 5-$n$ 90.12 $\pm 6.21$ 20-$n$ 87.05 $\pm 7.46$ Refer to caption
@@
In addition to max, we study the effect of using $m$-nearest neighbor (m-nn) as an aggregation method. Recall that max is a special case of m-nn when $m=1$ (as NORMpost is set to L2). We investigate different values for $m\in\{1,3,5,20,50\}$.
When all data is imprinted, fig.˜8 shows that using m-nn aggregation for $m\in\{3,5,20\}$ is slightly better than max. With $k=20$ and k-means as GEN, max (==1-$n$) aggregation becomes the top performing combination (see fig.˜9). Furthermore, the reduction of proxies (from all ($\approx 6000$ per class) to $k=20$) leads to a decrease in accuracy of less than $3\%$.
Figure 10: k-means with different values for $k$ in n-shot scenarios. Shaded areas indicate $95\%$ confidence intervals (CIs) over five different seeds. Other variables are fixed to our previously described best method of using L2 normalizations and max as AGG. Note that only data for the meaningful case of $k\leq n$ is shown. For MNIST and FashionMNIST, mean ceases to be the best strategy already at only roughly 50 samples.
5.2 Low-Data Regime
We analyze the n-shot scenario with our method (k-means as GEN, L2 as NORM, and max as AGG). Furthermore, we focus on the large tasks T containing all ten classes of MNIST, FashionMNIST, resp. CIFAR-10 at once. In this scenario, due to sampling only a few examples $n$, we average over five (instead of three) different seeds.
From the results in fig.˜10, we find that as the number of samples $n$ increases, k-means starts to outperform mean imprinting. The use of more proxies $k$ further amplifies this performance gain. The shift occurs at roughly 50 samples per class for MNIST and FashionMNIST, while for CIFAR-10, $k>1$ becomes prominently better at around 200 samples per class (see fig.˜A.4 for a display of all FMs focused on $10\leq n\leq 400$).
@@ 5.3 Neural Collapse and Number of Proxies
Figure˜11(a) depicts the neural collapse measurement $\mathcal{NC}_{1}$ (see eq.˜1) for the 100 random ImageNet tasks with remapped labels as explained in section˜4, as well as the four tasks containing all of MNIST, FashionMNIST, CIFAR-10, resp. CombiDigits. It can be inferred that ImageNet has a close-to-zero $\mathcal{NC}_{1}$ score, which increases linearly when adding more classes to each label (i.e., increasing multi-modality). As for the other datasets, CIFAR-10 is generally more collapsed according to its low value of $\mathcal{NC}_{1}$, which even falls below $1$ for the Transformer-based FMs. We hypothesize that this is due to the similarity of its categories to those appearing in ImageNet. By design, the synthetic CombiDigits dataset, introduced in section˜4, has a very high $\mathcal{NC}_{1}$. Apart from that, $\mathcal{NC}_{1}$ of the ImageNet data for the Transformer-based architectures are much lower and therefore they are more collapsed compared to the CNN-based FMs. These architectural differences are further investigated in section˜A.4.
(a) A clear linear relationship between the neural collapse measure $\mathcal{NC}_{1}$ and $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse. The $\mathcal{NC}_{1}$ of other datasets (CombiDigits in particular) is much higher across all FMs, while only vit_b_16 and swin_b get an $\mathcal{NC}_{1}$ of less than one on CIFAR-10.
@@
(b) Accuracy over number of proxies $k$ used in k-means together with L2 for all NORM and max as aggregation. In all four subplots, peaks in accuracy at $k=d$ can be inferred. This confirms the connection between the effect of using multiple proxies and the collapse of the data.
Figure 11: Averaged $\mathcal{NC}_{1}$ resp. accuracy of ten random ImageNet label remappings ($d$ in 11) for every $d=1,\dotsc,10$. Shaded areas indicate $95\%$ CIs over three different seeds. Values for the tasks containing all of MNIST, FashionMNIST, CIFAR-10, resp. CombiDigits at once are shown in dotted styles.
For the same data, fig.˜11(b) depicts accuracy over a varying number of proxies $k$ inferred from k-means. A prominent peak at $k=d$ can be inferred for every FM and reflects that $d$ class proxies lead to the best result for $d$-modal class distributions. Furthermore, increasing $k$ for the ImageNet sets has a much larger effect on the CNN-based FMs. We argue that this is because of their higher values of $\mathcal{NC}_{1}$, indicating less neural collapse. In section˜A.4, we analyze architectural differences and training setups to explain these variations. The fact that CIFAR-10 has the lowest $\mathcal{NC}_{1}$ scores (see fig.˜11(a)) is reflected by flat green curves over $k$.
Figure˜12 renders this connection more precise: For datasets with large $\mathcal{NC}_{1}$, i.e., where intra-class variability surpasses inter-class variability, using more than a single proxy per class ($k>1$) yields a clear performance gain over mean imprinting ($k=1$).
Figure 12: Accuracy (%) gain when using $k=5$ instead of $k=1$ proxies per class with k-means, plotted against the neural collapse measure $\mathcal{NC}_{1}$ (log-scaled axis). Colors indicate different FMs. Outlined markers correspond to non-ImageNet datasets (MNIST, FashionMNIST, CIFAR-10, and CombiDigits), while filled markers denote averaged results from our random ImageNet label remappings. The dashed gray line shows a least-squares fit ($R^{2}=0.36$), summarizing an approximate log-linear dependence between $\mathcal{NC}_{1}$ and the gains from using $k>1$ proxies. Across all points, the Spearman rank correlation is $\rho=0.82$ ($p<0.0001$), indicating a statistically significant positive association. Consistent with this trend, for regimes with $\mathcal{NC}_{1}>1$ single-proxy mean imprinting ($k=1$) appears to become substantially suboptimal compared to using multiple proxies ($k>1$).
6 Conclusion